OpenStudy (smurfy14):

how do i get rid of fractions in 9/7x+y+17/7=0? P.S.~keep it in general form

OpenStudy (anonymous):

I am assuming (9/7)x+y+17/7=0 Just multiply through by 7 to get 9x+7y+17=0

OpenStudy (anonymous):

9x+7y+17=0

OpenStudy (smurfy14):

alright thanks im assuming i did the work right because the original problem was to find the euation in general form for a line passing through (-5, 4) and (-2, 5)

OpenStudy (anonymous):

slope of these two points is 1/3

OpenStudy (smurfy14):

oh sorry it should say (-5,4) and (2, -5)

OpenStudy (anonymous):

never mind